# Compositional Difference-in-Differences for categorical outcomes

### v1

@@ 2.1. Relaxations of parallel growths
When multiple pre-treatment periods are available, the plausibility of the parallel growths assumption can be evaluated by examining the pre-treatment trajectories of the log-quantities. If the estimated pre-treatment trends appear non-parallel, this raises concerns about the validity of the assumption. Rather than discarding the design altogether, one can relax the assumption in a way that delivers informative bounds on the treatment effect. In particular, we adopt a relaxation strategy closely related to the approach in Ban and Kédagni (2022). They relax the parallel trends assumption by allowing the counterfactual difference between treated and control units to lie within the convex hull of the differences observed in the two most recent pre-treatment periods ($t=0$ and t=1t=-1). By taking the convex hull of the last two differences, they are saying: the counterfactual evolution cannot jump outside the range spanned by recent dynamics. If the differences between groups have been changing slowly, then the most recent differences provide a plausible envelope for what would have happened in the absence of treatment. This is weaker than requiring parallelism and rules out wild deviations inconsistent with observed history. Analogously, in our setting, we extend this idea by applying the same type of convex-hull relaxation. We consider three main relaxations, where we can either only use the two most recent periods, all the past periods, or all the past periods with higher weights to the most recent ones. In all the cases, we provide the corresponding bounds. We keep the same notation as before, but now, we have that $t\in\{-T_{1},...,0,1\}$. The three relaxations are given by the following assumptions:
@@
$\log\left(q_{1,1}^{N}(c_{k})\right)-\log\left({q_{0,1}^{N}(c_{k})}\right)\in\operatorname{Conv}\Bigg\{\log\left(\frac{q_{1,0}^{N}(c_{k})}{q_{0,0}^{N}(c_{k})}\right),\log\left(\frac{q_{1,-1}^{N}(c_{k})}{q_{0,-1}^{N}(c_{k})}\right)\Bigg\}.$
Assumption 4.
(Relaxation II)
For every category $c_{k}$ where $k=1,\dots,p$, we have:
$\log\left(q_{1,1}^{N}(c_{k})\right)-\log\left({q_{0,1}^{N}(c_{k})}\right)\in\operatorname{Conv}\Bigg\{\min_{t<0}\log\left(\frac{q_{1,t}^{N}(c_{k})}{q_{0,t}^{N}(c_{k})}\right),\max_{t<0}\log\left(\frac{q_{1,t}^{N}(c_{k})}{q_{0,t}^{N}(c_{k})}\right)\Bigg\}.$
Assumption 5.
(Relaxation III)
For every category $c_{k}$ where $k=1,\dots,p$, we have for $\omega_{t}\geq 0,\omega_{t}>\omega_{t-1}$
$\log\left(q_{1,1}^{N}(c_{k})\right)-\log\left({q_{0,1}^{N}(c_{k})}\right)\in\operatorname{Conv}\Bigg\{\min_{t<0}\omega_{t}\log\left(\frac{q_{1,t}^{N}(c_{k})}{q_{0,t}^{N}(c_{k})}\right),\max_{t<0}\omega_{t}\log\left(\frac{q_{1,t}^{N}(c_{k})}{q_{0,t}^{N}(c_{k})}\right)\Bigg\}.$
In all the cases, the assumption ensures that deviations from parallelism in the pre-treatment periods are accounted for, while still preserving enough structure to obtain informative bounds on the counterfactual. In practice, the researcher can consider only one of the case. Under Assumption 3, identification relies on a simple two-period comparison (the two most recent periods). Assumption 4 extends this logic to multiple pre-treatment periods, enabling the use of more temporal information, which may lead to wider bounds. Assumption 5 further introduces weighted relaxations through the parameters $\omega_{t}$. For instance, periods closer to the treatment date might receive larger weights if they are believed to be more informative about counterfactual dynamics. The bounds for the quantities would also translate into bounds for the distributions. The key observation to derive the bounds for the counterfactual is that the convex hull of any two vectors in $\mathbb{R}^{p}$ can be characterized componentwise by taking the minimum and maximum of each coordinate. Intuitively, this is because the convex hull is the smallest set containing both vectors and every possible weighted average of them. For each coordinate, all convex combinations lie between the smallest and largest value of that coordinate, so the lower and upper bounds are just the componentwise minimum and maximum. The identified set is the collection of all quantities that are consistent with the data and the model assumptions.
@@
$b_{min}^{k}=\exp\left(\min\left\{\log\left(\frac{q_{1,0}^{N}(c_{k})}{q_{0,0}^{N}(c_{k})}\right),\log\left(\frac{q_{1,-1}^{N}(c_{k})}{q_{0,-1}^{N}(c_{k})}\right)\right\}+\log(q_{0,1}^{N}(c_{k}))\right),$
$b_{max}^{k}=\exp\left(\max\left\{\log\left(\frac{q_{1,0}^{N}(c_{k})}{q_{0,0}^{N}(c_{k})}\right),\log\left(\frac{q_{1,-1}^{N}(c_{k})}{q_{0,-1}^{N}(c_{k})}\right)\right\}+\log(q_{0,1}^{N}(c_{k}))\right),$
Under assumptions 1, and 4, for all $\{c_{k},k=1,...,p$}, we have:
$b_{min}^{k}=\exp\left(\min\left\{\min_{t<0}\log\left(\frac{q_{1,t}^{N}(c_{k})}{q_{0,t}^{N}(c_{k})}\right),\min_{t<0}\log\left(\frac{q_{1,t}^{N}(c_{k})}{q_{0,t}^{N}(c_{k})}\right)\right\}+\log(q_{0,1}^{N}(c_{k}))\right),$
$b_{max}^{k}=\exp\left(\max\left\{\min_{t<0}\log\left(\frac{q_{1,t}^{N}(c_{k})}{q_{0,t}^{N}(c_{k})}\right),\min_{t<0}\log\left(\frac{q_{1,t}^{N}(c_{k})}{q_{0,t}^{N}(c_{k})}\right)\right\}+\log(q_{0,1}^{N}(c_{k}))\right),$
Under assumptions 1, and 5, for all $\{c_{k},k=1,...,p$}, we have:
$b_{min}^{k}=\exp\left(\min\left\{\min_{t<0}\omega_{t}\log\left(\frac{q_{1,t}^{N}(c_{k})}{q_{0,t}^{N}(c_{k})}\right),\min_{t<0}\omega_{t}\log\left(\frac{q_{1,t}^{N}(c_{k})}{q_{0,t}^{N}(c_{k})}\right)\right\}+\log(q_{0,1}^{N}(c_{k}))\right),$
$b_{max}^{k}=\exp\left(\max\left\{\min_{t<0}\omega_{t}\log\left(\frac{q_{1,t}^{N}(c_{k})}{q_{0,t}^{N}(c_{k})}\right),\min_{t<0}\omega_{t}\log\left(\frac{q_{1,t}^{N}(c_{k})}{q_{0,t}^{N}(c_{k})}\right)\right\}+\log(q_{0,1}^{N}(c_{k}))\right),$
In all the cases, we have the following bounds for the treatment effect:
$GTT(c_{k})\in\left[\frac{q_{1,1}^{I}(c_{k})}{b_{max}^{k}}-1,\frac{q_{1,1}^{I}(c_{k})}{b_{min}^{k}}-1\right],$
$GTT\in\left[\frac{\sum_{k=1}^{p}q_{1,1}^{I}(c_{k})}{\sum_{k=1}^{p}b_{max}^{k}}-1,\frac{\sum_{k=1}^{p}q_{1,1}^{I}(c_{k})}{\sum_{k=1}^{p}b_{max}^{k}}-1,\right]$
$CTT\in\left\{\pi\in\mathcal{S}^{p-1}:\pi=\pi^{I}_{1,1}\ominus s,s\in\ell^{-1}\left(\left\{r\in\mathbb{R}^{p-1}\,\Big|\,\log(b_{min}^{k})\leq\ r_{k}\leq\log(b_{max}^{k})\right\}\right)\right\}.$
Theorem 2 provides a unified set of partial identification results for the counterfactual treatment effects under progressively different relaxations. Each set of assumptions 3, 4, and 5 represents a distinct way to relax the canonical parallel growths assumption, allowing for more flexibility while maintaining interpretability and structure in the evolution of shares across treatment and control groups. In all three cases, the bounds $b_{\min}^{k},b_{\max}^{k}]$ define the identified set for the counterfactual category-level quantities, which in turn determine bounds for three related treatment effect measures.
@@ 2.2. Staggered adoption
In this section, we extend the staggered treatment adoption framework of Callaway and Sant’Anna (2021), originally developed for scalar outcomes, to the setting of categorical outcomes. Such designs are increasingly common in applied policy evaluation, where reforms or programs are rolled out gradually across units rather than implemented simultaneously. We consider a panel of $n$ units observed over periods $t=0,1,\dots,T$. Let $D_{i,t}\in\{0,1\}$ indicate whether unit $i$ is treated in period $t$. We impose the standard irreversibility assumption:
@@
Under this assumption, each unit that ever receives treatment has a well-defined first treatment time, denoted $G_{i}=\min\{t:D_{i,t}=1\}$. Units never treated are assigned $G_{i}=\infty$. This partitions the sample into cohorts: for each $g\in\{1,\dots,T\}$, group $G_{g}$ consists of units first treated in period $g$; the never-treated group is $G_{\infty}$. Our outcome of interest is categorical: for each group and time, we observe a vector of positive quantities $q_{G_{g},t}$. We also work with the associated share vector $\pi_{G_{g},t}=q_{G_{g},t}/\sum_{k=1}^{p}q_{G_{g},t}(c_{k})$.
For each cohort $G_{g}$, we observe pre-treatment (untreated) outcomes for all $t<g$, and post-treatment (treated) outcomes for $t\geq g$. Our goal is to identify the counterfactual untreated quantity $q_{G_{g},t}^{N}$ for $t\geq g$. Let $\bar{g}=\max\{G_{i}:G_{i}<\infty\}$ denote the last period in which any unit is first treated, and define $\mathcal{G}=\text{Supp}(G)\setminus\{\bar{g}\}$ as the set of all treated cohorts except the last one. (The last cohort cannot be used as a control for itself under our second identification strategy, as explained below). We propose two identification strategies, both generalizing the parallel growths assumption (assumption 2) to the staggered setting:
@@ 2.3. Synthetic CoDiD
Now suppose we have a setting with $G>2$ groups observed over $T>2$ time periods. Only the first group ($g=1$) receives treatment, beginning after period $T_{0}$; all other groups ($g=2,\dots,G$) serve as never-treated controls. In this context, Arkhangelsky et al. (2021) propose the Synthetic difference-in-differences (SDID) method, an extension of the classic DiD framework that constructs a data-driven synthetic control by optimally reweighting the control units to best reproduce the pre-treatment trajectory of the treated unit. This improves the credibility of post-treatment comparisons by enforcing a stronger pre-treatment fit than standard DiD. In this section, we adapt SDID to settings where the outcome is a categorical quantity vector—for example, expenditures across categories, counts of events by type, or any positive-valued composition. For each group $g$ and time $t$, let $q_{g,t}$ denote the vector of quantities across $p$ categories, and let $T_{0}$ be the last pre-treatment period. Our objective is to estimate treatment effects on both GTT and CTT. The procedure is applied category by category; for clarity of exposition, we suppress the category index $(c_{k})$ in what follows, noting that all steps are repeated for each component of the quantity vector.
  • Find Unit Weights($\boldsymbol{\omega}^{\ast}$): We find weights for the control groups so that their weighted pre-treatment path best matches the pre-treatment path of the treated group.
@@
  • Find Time Weights ($\boldsymbol{\lambda}^{\ast}$): We find weights for the pre-treatment time periods so that the pre-treatment average of the controls best matches their post-treatment average.
@@
Finally, we calculate the average growth treatment effect.
@@
We then construct four key average distributions by normalizing across categories, we obtain the corresponding distributions: $\boldsymbol{\pi}_{\text{treated,pre}}$, $\boldsymbol{\pi}_{\text{treated,post}}$, $\boldsymbol{\pi}_{\text{control,pre}}$, $\boldsymbol{\pi}_{\text{control,post}}$.
In our probability space, we use operations $\ominus$ (perturbation-difference) and $\oplus$ (perturbation-addition) instead of regular subtraction and addition. The Compositional Treatment Effect (CTT) is defined as:
@@ 3. Connexions With Existing Approaches
In this section, we discuss the link between our method and existing methods in the literature.
@@ 3.1. Connexion with Logistic model
This specification is fully saturated: for each non-baseline outcome category $c_{k}$, the model includes four parameters, an intercept $\alpha_{k}$, a group effect $\delta_{k}$, a time effect $\gamma_{k}$, and an interaction term $\beta_{k}$, which together exactly span the log-odds in all four group–time cells. The interaction coefficient $\beta_{k}$ captures the difference-in-differences contrast in log-odds and admits a causal interpretation under the assumption of parallel trends on the log-odds. Our compositional treatment effect parameter (CTT) can be linked to the coefficients of a multinomial logit model as follows:
@@
where $\ell^{-1}(\cdot)$ denotes the inverse compositional log-ratio transformation, which maps the vector of log-odds coefficients back into category shares. Although using the multinomial logit representation coincides with our approach to modeling category shares in the simple $(2\times 2)$ setup, several conceptual and practical differences are worth emphasizing.
The first key difference is that the multinomial logit model does not directly yield counterfactual quantities without imposing additional structural assumptions. In contrast, our approach provides these counterfactuals constructively, enabling the recovery of both treatment effects on quantities and on shares within a unified framework. Second, in settings with multiple time periods, the multinomial logit specification offers no straightforward way to relax the identifying assumptions when they appear violated. Our framework, however, accommodates such relaxations naturally, formalized through assumptions 3, 4, and 5 allowing for partial identification. Third, when treatment timing varies across units (staggered adoption), the multinomial logit model requires estimating a full vector of time-varying treatment effect coefficients. By contrast, our method allows the estimation of treatment effect for each post-treatment period independently, without the need for joint optimization or functional minimization. This feature makes our approach computationally lighter. Finally, the Compositional Treatment Effect (CTT) parameter provides a direct and interpretable measure of the treatment’s impact on category shares. Operating within the space of compositional data, it connects naturally to discrete choice theory and can be naturally used in applications where proportions or probabilities, rather than levels, constitute the primary outcomes of interest.
@@ 3.2. Unified view with DiD and CiC
Our method identifies the counterfactual probability mass function for unordered categorical outcomes. We show how the identifying assumption relates to parallel trends (DiD) and rank invariance (CiC). DiD targets objects like expectations, CiC targets the CDF for ordered outcomes, and CoDiD extends the same logic to the PMF for unordered outcomes. We proceed in two steps. First, we show that DiD and CiC, although targeting different objects (means versus CDFs), share the same geometric identification logic. Second, we show how CoDiD applies that logic to probability mass functions. A recall of the setup: we consider two groups $g\in{0,1}$ (treatment and control), observed in two periods $t\in{0,1}$. Let $m_{g,t}^{N}$ denote the potential parameter of interest for unit $g$ at time $t$ in the absence of treatment.
@@
DiD and CiC share a simple underlying idea: the object of interest exists within a space where the notion of movement can be well-defined, and there is a transformation that characterizes how the control group evolves within this space. To obtain the counterfactual for the treated group, we then apply the same transformation to the treated group. Mathematically, if the control group moves from $m_{0,0}^{N}$ to $m_{0,1}^{N}$, we write:
@@
where $\mathcal{F}$ is the space of cumulative distribution functions. The map $\varphi^{c}$ has the same interpretation in both methods and can be understood as a translation map. If we want to use a map to define the idea of movement within a space. A natural way to formalize that is through the concept of translation. A translation operator is defined as an operator that shifts points by a certain amount in a certain direction. We consider a general definition of this operator. Intuitively, suppose the outcome space forms a group, meaning it is a set of objects equipped with a binary operation that can help define the idea of movement (moves that can be combined, reversed, and have a neutral element that does nothing). These moves can be addition (for numeric outcomes) or function composition (for distributional transformations). Let’s start with the formal definition of a group.
@@
For example, in DiD, the outcome space is $\mathbb{R}$ with addition ++ as the group operation, identity $e=0$, and inverse given by negation a-a. In CiC, the outcome space is the set of cumulative distribution functions, the operation is composition $\circ$, the identity is the identity function $Id(x)=x$, and the inverse is the quantile function. Once the group structure is defined, we can formalize the idea of a movement using translations.
@@
If you start at $m$, you then apply $h$ on the right, ending at $m\wedge h$. Think of shifting relative to where you already are, like taking your current position and then performing the move locally. This formalizes the notion of a shift or transformation acting consistently across the entire outcome space. From this perspective, the map $\varphi^{c}$ can be identified as the unique right translation that captures how the control group evolves over time.
@@
Graphically, we can illustrate through figure 5.
$\mathcal{W}$$F_{Y^{N}_{0,0}}$$F_{Y^{N}_{0,1}}$$F_{Y^{N}_{1,0}}$$F_{Y^{N}_{1,1}}$
@@
That unified logic breaks down if we use the Euclidean group structure for the simplex (doing parallel trends on proportions). In fact, when we restrict to the simplex, the structure no longer forms a group since the neutral element would have to be the zero vector (which is not part of the simplex). In contrast, under our formulation, the group structure is preserved, ensuring consistency with DiD and CiC. Table 1 summarizes this comparison. Let $\mathcal{F}$, be the space of non-decreasing real functions.

### v2

@@ 2.1. Relaxations of parallel growths
When multiple pre-treatment periods are available, the plausibility of the parallel growths assumption can be evaluated by examining the pre-treatment trajectories of the log-quantities. If the estimated pre-treatment trends appear non-parallel, this raises concerns about the validity of the assumption. Rather than discarding the design altogether, one can relax the assumption in a way that delivers informative bounds on the treatment effect. In particular, I adopt a relaxation strategy closely related to the approach in Ban and Kédagni (2022). They relax the parallel trends assumption by allowing the counterfactual difference between treated and control units to lie within the convex hull of the differences observed in the two most recent pre-treatment periods ($t=0$ and t=1t=-1). By taking the convex hull of the last two differences, they are saying: the counterfactual evolution cannot jump outside the range spanned by recent dynamics. If the differences between groups have been changing slowly, then the most recent differences provide a plausible envelope for what would have happened in the absence of treatment. This is weaker than requiring parallelism and rules out wild deviations inconsistent with observed history. Analogously, in our setting, I extend this idea by applying the same type of convex-hull relaxation. I consider a more general case, where one can either only use the two most recent periods, all the past periods, or all the past periods with higher weights to the most recent ones. I keep the same notation as before, but now, we have that $t\in\{-T_{1},...,0,1\}$.
@@
$\log\left(q_{1,1}^{N}(c_{k})\right)-\log\left({q_{0,1}^{N}(c_{k})}\right)\Bigg[\min_{t\leq 0}\quad\omega_{t}\log\left(\frac{q_{1,t}^{N}(c_{k})}{q_{0,t}^{N}(c_{k})}\right),\quad\max_{t\leq 0}\quad\omega_{t}\log\left(\frac{q_{1,t}^{N}(c_{k})}{q_{0,t}^{N}(c_{k})}\right)\Bigg].$
[…]
@@
$GTT\in\left[\frac{\sum_{k=1}^{p}q_{1,1}^{I}(c_{k})}{\sum_{k=1}^{p}b_{max}(c_{k})}-1,\frac{\sum_{k=1}^{p}q_{1,1}^{I}(c_{k})}{\sum_{k=1}^{p}b_{min}(c_{k})}-1\right],$
$CTT\in\left\{\pi\in\mathcal{S}^{p-1}:\pi=\pi^{I}_{1,1}\ominus s,s\in\ell^{-1}\left(\left\{r\in\mathbb{R}^{p-1}\,\Big|\,\log(b_{min}(c_{k}))\leq\ r_{k}\leq\log(b_{max}(c_{k})),\right\}\right)\right\}$
where we have:
$b_{min}(c_{k})=\exp\left(\min_{t\leq 0}\omega_{t}\log\left(\frac{q_{1,t}^{N}(c_{k})}{q_{0,t}^{N}(c_{k})}\right)+\log(q_{0,1}^{N}(c_{k}))\right),$
$b_{max}(c_{k})=\exp\left(\max_{t\leq 0}\omega_{t}\log\left(\frac{q_{1,t}^{N}(c_{k})}{q_{0,t}^{N}(c_{k})}\right)+\log(q_{0,1}^{N}(c_{k}))\right).$
Theorem 2 provides a clear characterization of how partial identification operates under the relaxed parallel growths assumption. It establishes sharp bounds on both the growth treatment effect on the treated and the compositional treatment effect on the treated by allowing for deviations from parallelism between treated and control groups. The CTT bounds describe a feasible set within the simplex of compositions, capturing all possible redistributions consistent with those counterfactuals. The choice of $\omega_{t}$ governs the degree of relaxation: uniform weights yield a fully agnostic specification based solely on pre-treatment extremes, while time-increasing weights change the bounds by prioritizing near-treatment dynamics.
@@ 2.2. Staggered adoption
In this section, I extend the staggered treatment adoption framework of Callaway and Sant’Anna (2021), originally developed for scalar outcomes, to the setting of categorical outcomes. Such designs are increasingly common in applied policy evaluation, where reforms or programs are rolled out gradually across units rather than implemented simultaneously. I consider a panel of $n$ units observed over periods $t=0,1,\dots,T$. Let $D_{i,t}\in\{0,1\}$ indicate whether unit $i$ is treated in period $t$. I impose the standard irreversibility assumption:
@@
Under this assumption, each unit that ever receives treatment has a well-defined first treatment time, denoted $G_{i}=\min\{t:D_{i,t}=1\}$. Units never treated are assigned $G_{i}=\infty$. This partitions the sample into cohorts: for each $g\in\{1,\dots,T\}$, group $G_{g}$ consists of units first treated in period $g$; the never-treated group is $G_{\infty}$. The outcome of interest is categorical: for each group and time, I observe a vector of positive quantities $q_{G_{g},t}$. I also work with the associated share vector $\pi_{G_{g},t}=q_{G_{g},t}/\sum_{k=1}^{p}q_{G_{g},t}(c_{k})$.
For each cohort $G_{g}$, I observe pre-treatment (untreated) outcomes for all $t<g$, and post-treatment (treated) outcomes for $t\geq g$. The goal is to identify the counterfactual untreated quantity $q_{G_{g},t}^{N}$ for $t\geq g$. Let $\bar{g}=\max\{G_{i}:G_{i}<\infty\}$ denote the last period in which any unit is first treated, and define $\mathcal{G}=\text{Supp}(G)\setminus\{\bar{g}\}$ as the set of all treated cohorts except the last one. (The last cohort cannot be used as a control for itself under the second identification strategy, as explained below). I propose two identification strategies, both generalizing the parallel growths assumption (assumption 2) to the staggered setting:
@@ 2.3. Synthetic CoDiD
Now suppose we have a setting with $G>2$ groups observed over $T>2$ time periods. Only the first group ($g=1$) receives treatment, beginning after period $T_{0}$; all other groups ($g=2,\dots,G$) serve as never-treated controls. In this context, Arkhangelsky et al. (2021) propose the Synthetic difference-in-differences (SDID) method, an extension of the classic DiD framework that constructs a data-driven synthetic control by optimally reweighting the control units to best reproduce the pre-treatment trajectory of the treated unit. This improves the credibility of post-treatment comparisons by enforcing a stronger pre-treatment fit than standard DiD. In this section, I adapt SDID to settings where the outcome is a categorical quantity vector—for example, expenditures across categories, counts of events by type, or any positive-valued composition. For each group $g$ and time $t$, let $q_{g,t}$ denote the vector of quantities across $p$ categories, and let $T_{0}$ be the last pre-treatment period. The objective is to estimate treatment effects on both GTT and CTT. The procedure is applied category by category; for clarity of exposition, I suppress the category index $(c_{k})$ in what follows, noting that all steps are repeated for each component of the quantity vector.
  • Find Unit Weights($\boldsymbol{\omega}^{\ast}$): I find weights for the control groups so that their weighted pre-treatment path best matches the pre-treatment path of the treated group.
@@
  • Find Time Weights ($\boldsymbol{\lambda}^{\ast}$): I find weights for the pre-treatment time periods so that the pre-treatment average of the controls best matches their post-treatment average.
@@
Finally, I calculate the average growth treatment effect.
@@
We then construct the distributions by normalizing across categories: $\boldsymbol{\pi}_{\text{treated,pre}}$, $\boldsymbol{\pi}_{\text{treated,post}}$, $\boldsymbol{\pi}_{\text{control,pre}}$, $\boldsymbol{\pi}_{\text{control,post}}$.
The Compositional Treatment Effect (CTT) is defined as:
@@ 3. Connexions With Existing Approaches
In this section, I discuss the link between my method and existing methods in the literature.
@@ 3.1. Connexion with Logistic model
This specification is fully saturated: for each non-baseline outcome category $c_{k}$, the model includes four parameters, an intercept $\alpha_{k}$, a group effect $\delta_{k}$, a time effect $\gamma_{k}$, and an interaction term $\beta_{k}$, which together exactly span the log-odds in all four group–time cells. The interaction coefficient $\beta_{k}$ captures the difference-in-differences contrast in log-odds and admits a causal interpretation under the assumption of parallel trends on the log-odds. The compositional treatment effect parameter (CTT) can be linked to the coefficients of a multinomial logit model as follows:
@@
where $\ell^{-1}(\cdot)$ denotes the inverse compositional log-ratio transformation, which maps the vector of log-odds coefficients back into category shares. Although using the multinomial logit representation coincides with my approach to modeling category shares in the simple $(2\times 2)$ setup, several conceptual and practical differences are worth emphasizing.
The first key difference is that the multinomial logit model does not directly yield counterfactual quantities without imposing additional structural assumptions. In contrast, my approach provides these counterfactuals constructively, enabling the recovery of both treatment effects on quantities and on shares within a unified framework. Second, in settings with multiple time periods, the multinomial logit specification offers no straightforward way to relax the identifying assumptions when they appear violated. My framework, however, accommodates such relaxations naturally, formalized through assumptions 3 allowing for partial identification. Third, when treatment timing varies across units (staggered adoption), the multinomial logit model requires estimating a full vector of time-varying treatment effect coefficients. By contrast, my method allows the estimation of treatment effect for each post-treatment period independently, without the need for joint optimization or functional minimization. This feature makes my approach computationally lighter. Finally, the Compositional Treatment Effect (CTT) parameter provides a direct and interpretable measure of the treatment’s impact on category shares. Operating within the space of compositional data, it connects naturally to discrete choice theory and can be naturally used in applications where proportions or probabilities, rather than levels, constitute the primary outcomes of interest.
@@ 3.2. Unified view with DiD and CiC
My method identifies the counterfactual probability mass function for unordered categorical outcomes. I show how the identifying assumption relates to parallel trends (DiD) and rank invariance (CiC). DiD targets objects like expectations, CiC targets the CDF for ordered outcomes, and CoDiD extends the same logic to the PMF for unordered outcomes. I proceed in two steps. First, I show that DiD and CiC, although targeting different objects (means versus CDFs), share the same geometric identification logic. Second, I show how CoDiD applies that logic to probability mass functions. A recall of the setup: I consider two groups $g\in{0,1}$ (treatment and control), observed in two periods $t\in{0,1}$. Let $m_{g,t}^{N}$ denote the potential parameter of interest for unit $g$ at time $t$ in the absence of treatment.
@@
DiD and CiC share a simple underlying idea: the object of interest exists within a space where the notion of movement can be well-defined, and there is a transformation that characterizes how the control group evolves within this space. To obtain the counterfactual for the treated group, we then apply the same transformation to the treated group. Mathematically, if the control group moves from $m_{0,0}^{N}$ to $m_{0,1}^{N}$, I write:
@@
where $\mathcal{F}$ is the space of cumulative distribution functions. The map $\varphi^{c}$ has the same interpretation in both methods and can be understood as a translation map. If I want to use a map to define the idea of movement within a space. A natural way to formalize that is through the concept of translation. A translation operator is defined as an operator that shifts points by a certain amount in a certain direction. I consider a general definition of this operator. Intuitively, suppose the outcome space forms a group, meaning it is a set of objects equipped with a binary operation that can help define the idea of movement (moves that can be combined, reversed, and have a neutral element that does nothing). These moves can be addition (for numeric outcomes) or function composition (for distributional transformations). Let’s start with the formal definition of a group.
@@
For example, in DiD, the outcome space is $\mathbb{R}$ with addition ++ as the group operation, identity $e=0$, and inverse given by negation a-a. In CiC, the outcome space is the set of cumulative distribution functions, the operation is composition $\circ$, the identity is the identity function $Id(x)=x$, and the inverse is the quantile function. Once the group structure is defined, I can formalize the idea of a movement using translations.
@@
If you start at $m$, you then apply $h$ on the right, ending at $m\wedge h$. Think of shifting relative to where you already are, like taking the current position and then performing the move locally. This formalizes the notion of a shift or transformation acting consistently across the entire outcome space. From this perspective, the map $\varphi^{c}$ can be identified as the unique right translation that captures how the control group evolves over time.
@@
Graphically, I can illustrate through figure 5.
$\mathcal{F}$$F_{Y^{N}_{0,0}}$$F_{Y^{N}_{0,1}}$$F_{Y^{N}_{1,0}}$$F_{Y^{N}_{1,1}}$
@@
That unified logic breaks down if I use the Euclidean group structure for the simplex (doing parallel trends on proportions). In fact, when I restrict to the simplex, the structure no longer forms a group since the neutral element would have to be the zero vector (which is not part of the simplex). In contrast, under my formulation, the group structure is preserved, ensuring consistency with DiD and CiC. Table 1 summarizes this comparison. Let $\mathcal{F}$, be the space of non-decreasing real functions.